# Evaluating from Benign to Dynamic Adversarial: A Squid Game for Large Language Models

### v1

@@ Abstract
Contemporary benchmarks are struggling to keep pace with the development of large language models (LLMs). Although they are indispensable to evaluate model performance on various tasks, it is uncertain whether the models trained on Internet data have genuinely learned how to solve problems or merely seen the questions before. This potential data contamination issue presents a fundamental challenge to establishing trustworthy evaluation frameworks. Meanwhile, existing benchmarks predominantly assume benign, resource-rich settings, leaving the behavior of LLMs under pressure unexplored. In this paper, we introduce Squid Game, a dynamic and adversarial evaluation environment with resource-constrained and asymmetric information settings elaborated to evaluate LLMs through interactive gameplay against other LLM opponents. Notably, Squid Game consists of six elimination-style levels, focusing on multi-faceted abilities, such as instruction-following, code, reasoning, planning, and safety alignment. We evaluate over 50 LLMs on Squid Game, presenting the largest behavioral evaluation study of general LLMs on dynamic adversarial scenarios. We observe a clear generational phase transition on performance in the same model lineage and find evidence that some models resort to speculative shortcuts to win the game, indicating the possibility of higher-level evaluation paradigm contamination in static benchmarks. Furthermore, we compare prominent LLM benchmarks and Squid Game with correlation analyses, highlighting that dynamic evaluation can serve as a complementary part for static evaluations. The code and data will be released in the future.
@@ 1 Introduction
Evaluation has always been an important cornerstone for large language model (LLM) development, with a proliferation of general benchmarks (Bai et al., 2024; Wang et al., 2024; White et al., 2025; Chen et al., 2025b) and their extensions flourishing across multiple domains, covering specific disciplines (Sun et al., 2024; Chen et al., 2025a; Rein et al., 2024), reasoning Valmeekam et al. (2023), code (Jain et al., 2025), safety (Zhang et al., 2024b), and multi-modality (Li et al., 2024a; Fu et al., 2024; Yue et al., 2024; Chen et al., 2025c) tasks. Despite these rapid advancements, evaluation methodologies have remained relatively stagnant. Current benchmarks predominantly conduct evaluations under benign environments with abundant computational resources (e.g., no quota limit) and stable interaction (e.g., fixed question-answer format), resulting in a pronounced gap between their high theoretical performance and practical utility. Moreover, most of them are static, closed-ended, and composed of knowledge-intensive tasks, making them susceptible to data contamination from the pre-training corpus (Xu et al., 2024; Deng et al., 2024).
@@
Figure 1: Left. A simplified tournament bracket of the Squid Game, showing the advancement path of competing models through successive elimination rounds. Squid Game comprises levels from static parallel to dynamic adversarial settings, allowing all-round evaluations. Right. The figure depicts pass rates of o4-mini, Gemini 2.5 Pro, Grok-4, Seed 1.6, Kimi K2, and Qwen3-32B over different levels in 20 rounds of evaluation.
@@
Motivated by these shortcomings, we introduce Squid Game, built on the following principles:
@@
    Elimination rather than Score. Traditional score-based benchmarks (Fu et al., 2024; Wang et al., 2024; Yue et al., 2024; Chen et al., 2025a) conduct evaluations by instructing models to complete an examination independently in a static and resource-unlimited environment. Afterwards, an absolute score is obtained that represents the model’s mastery over a specific body of knowledge, essentially answering the question “How much do you know?”. Such a scoring system could lead to the phenomenon of ranking fraud, where model developers may over-optimize for the specific question types, constituting a form of exam-oriented training. In Squid Game, instead of seeking the theoretically best model, we adopt a Battle Royale-style relative ranking system, where a certain percentage of the models may be eliminated in each round. Based on this, the difficulty of following challenges increases according to both the game settings and the survivors themselves. Therefore, a high ranking not only represents superior capabilities but also reflects a tactical advantage (e.g., fewer errors, better strategies) over specific opponents. The final pass rate distribution demonstrated in Fig. 1 illustrates the difference between static evaluation results and dynamic evaluation results.
@@
    Resource Constraint. With the increasing scale and breadth of benchmarks, the demand for evaluation resources has also grown accordingly. Running thoroughly on MMMU-Pro (Yue et al., 2024) for a 72B LLM requires more than two GPUs, each with 80GB of VRAM, taking several dozen hours to complete. As for API models, evaluation costs can run into the tens of dollars, which is magnified by modern reasoning models that inflate costs by generating a large volume of verbose and often superfluous output tokens. The previous construction mode of LLM benchmarks encourages the development of larger models with more training data. However, does this mean they are more intelligent, or simply a larger database? To this end, we add the resource constraint (e.g., token or API call quota) in Squid Game to measure the utilization rate of tokens and inference efficiency. More importantly, this mechanism also exerts pressure on the model by providing live feedback on its diminishing resource pool.
@@
    Information Asymmetry. High-quality question designs and answer forms are crucial for reliable evaluation of LLMs. However, existing benchmarks suffer from multiple deficiencies. First, leading benchmarks, such as MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), MMMU (Yue et al., 2024), and LiveBench (White et al., 2025), have almost all adopted multi-choice, binary judgment, or open-ended forms under fixed question-answer (QA) scenarios, where the core assumption is one of complete information, i.e., the prompt itself provides all necessary context. Furthermore, option-based evaluation is more susceptible to hallucination and output preferences (Chen et al., 2025c), resulting in biased observations. In contrast, we introduce an information asymmetry design that shifts the evaluative focus from “what a model knows” to “what it can do under uncertainty”. For example, in glass stepping stones, the participating LLMs start with no prior information, unaware of whether any given glass pane is safe or a trap. The dynamic asymmetry between models lies in that those positioned further down the order hold a substantial informational advantage, as they can observe the actions and results of all prior competitors. This uncertainty of the evaluation process improves the depth of evaluation and can differentiate models with better reasoning, planning, and strategy generation capabilities towards artificial general intelligence (AGI).
  4. 4.
    Dynamic Adversarial Evaluation. Current benchmarks are predominantly static and closed-ended with independent evaluation processes, which makes them prone to data contamination, leading to untrustworthy, inflated results that do not accurately reflect their true capabilities (Magar & Schwartz, 2022; Xu et al., 2024; Deng et al., 2024). Recent works (Chan et al., 2024; Gao et al., 2025) have demonstrated the feasibility of the interactive and collaborative evaluation paradigm for LLMs. Google’s recent chess tournament for LLMs (Google, 2025) serves as a good example. In Squid Game, we elaborate a series of levels (e.g., red-green light, tug of war, and marbles) that provide a holistic evaluation of all-round capabilities, ranging from benign instruction following to collaborative problem-solving and adversarial gaming in both offensive and defensive scenarios. This creates a self-evolving and never-saturating evaluation environment, where the difficulty automatically scales with the opponent’s intelligence, thus maintaining a constant challenge for cutting-edge models.
With these principles in mind, we build Squid Game, a dynamic adversarial game-form benchmark that exposes LLMs to extreme stress-testing conditions. Inspired by the globally popular dystopian drama series, “Squid Game”, produced by Netflix, we construct six different evaluation scenarios.
Empirical Findings. We have evaluated over 50 LLMs (28 proprietary and 24 open-source) across different Squid Game scenarios. Based on this study and the following analysis, we present novel empirical findings that have not been revealed in prior benchmarks.
  1. 1.
    Holistic Evaluation. Our evaluations reveal that model dynamic performances is a multifaceted outcome, determined not only by task complexity but also by its architectural class (e.g., lightweight and inherent reasoning capabilities), rather than a monolithic scaling law.
  2. 2.
    Constructing Dynamic Evaluation. We observe a potential flaw on the traditional evaluation methodology, which is as important as data contamination. A paradigm shift from static benchmarks to dynamic evaluation frameworks is essential to overcome this limitation.
  3. 3.
@@
    Complementary to Static Evaluation. Low correlation is observed between the results and static benchmarks, indicating the necessity of exploring the dynamic abilities of LLMs in current landscape, since the performance on static benchmarks is approaching saturation.
Concurrent Work. The existing dynamic evaluation approaches primarily focus on the transformation of data. DyVal (Zhu et al., 2024) and DARG (Zhang et al., 2024a) leverage the graph structure of data points in current benchmarks and generate novel testing data. Some studies (Wang et al., 2025; Chen et al., 2024) utilize multi-agent systems to generate real-time variable instances or reframe new ones for self-evolving benchmarks. However, these approaches neglect the potential conceptual and methodological leakage that, rather than memorizing specific test instances, the model may have internalized the underlying solution patterns or heuristics for a given problem archetype from its training corpus. Recent Game Arena conducted by Kaggle and Google (Olszewska & Risdal, 2025) introduces a chess-only adversarial evaluation. Our Squid Game aims to overcome these hurdles by both introducing richer dynamic adversarial evaluation scenarios and neutralizing meta-knowledge leakage with information asymmetry, variable process, and adaptive difficulty level.
Figure 2: Overview of the six scenarios present in Squid Game.
@@ 2 Constructing the Squid Game
Games provide an intuitive and comparable signal of success. The randomness of the process and structured nature make them suitable for evaluating models and agents, which have been widely used since the era of AlphaGo. Unlike the unidirectional question-answer format-based evaluations, they compel models to demonstrate many capabilities, including strategic reasoning, planning, and survivability, under competitive scenarios, offering a stepwise visualization of their behavior. Such game-based behavioral research is also quite popular in other fields. Previous prevailing Netflix series, Squid Game, a South Korean survival drama with critical acclaim, has been widely studied in the sociology domain (Aoun, 2022). The high-pressure game settings serve to magnify each participant’s capabilities and human frailties, which satisfy the reform requirements of today’s LLM evaluation.
Motivated by this, we build a full-scale replica of the Squid Game for general LLMs in six scenarios, namely red-green light, sugar honeycombs, tug of war, marbles, glass stepping stones, and the final squid game, as shown in Fig. 2. Following we describe each of these scenarios in detail.
Red-Green Light. The goal of this game was for players to successfully reach the other end of an enclosed field within a certain time limit. Players can move on a green light but must freeze on a red light. During the red light scan, any player caught moving will be eliminated. In this scenario, we focus on the accuracy and timeliness of the instruction-following ability of LLMs. Considering that the standard API call or local execution patterns for mainstream LLMs do not permit the external injection of a new, real-time interrupt command during the inference process, we adopt a turn-based long task decomposition strategy:
  • Rule Definition: At the beginning of the game, the model receives a system prompt $P_{system}$ that defines the complete requirement of engagement. This includes the general description of the task and a rule for generating the security code.
@@
  • Game Loop: The game proceeds in turns. In each round, two types of commands are randomly generated, namely “Continue” (green light) for generating the next chunk of the long task and “Interruption” (red light) for immediately stopping the main task and outputting the security code according to the predefined rule. A verification function is deployed to eliminate the models that do not meet the code requirements. The model then takes the output from the previous round along with the next sub-task instruction and repeats the process iteratively.
The state of the $i$-th round can be formulated as:
$\left(X_{i}\right)_{i\in\mathbb{N}_{+}}=\begin{cases}I_{i}=P_{task},\ H_{i}=P_{system}\oplus I_{i}&i=1\\ I_{i+1}\sim\begin{cases}\text{"Interrupt"}&:\rho\\ \text{"Continue"}&:1-\rho\end{cases},\ H_{i+1}=H_{i}\oplus M\left(H_{i}\right)\oplus I_{i+1}&1\leq i\leq N\end{cases}$ (1)
where $P_{task}$ denotes the initial long task prompt. $I_{i}$ and $H_{i}$ are the instruction and the cumulative context history of $i$-th round. $M(\cdot)$ denotes the response of model. $\rho$, $N$, and $\oplus$ represent the probability of an interruption instruction, the number of rounds, and the string concatenation, respectively.
@@
Sugar Honeycombs. In this game, contestants were tasked with using a needle to carve a shape out of a piece of dalgona without cracking the honeycomb. Players who broke their honeycombs or failed to finish within the time limit were eliminated from the game. Based on this, we focus on the stability and precise control ability of models for performing delicate operations in demanding conditions. Specifically, we design a code sculpting scenario where the models are given a functionally complete but structurally suboptimal code $C_{in}$ with poor readability, high complexity, or redundant logic. This task is to refactor $C_{in}$ into a new version $C_{out}$ while satisfying a set of highly restrictive rules $\mathcal{R}$:
@@
where the survival function $\mathbb{I}_{survival}$ equals $1$ if the condition is met, and 0 if even a single test case fails. $\mathbb{I}(\cdot)$ is the indicator function and $C_{out}=M(C_{in},\mathcal{R})$. $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ includes the set of $n$ unit tests. This evaluation framework shifts the focus from what a model can create or fix to how well it can understand and precisely manipulate existing functional code.
@@
Tug of War. Compared to the first two levels that were accomplished independently and implemented in parallel, this game assesses the teamwork capabilities across multiple models and the overall effect of their collaboration under random allocation. This game involves two teams of three (i.e., the lead debater, supporting debater, and summarizer) debating for a given topic, where both sides engage in multiple rounds of language games, generating arguments, rebuttals, and attempting to steer the consensus of the central issue towards the position they had pre-determined (Estornell & Liu, 2024). A hybrid judging system, including a superior LLM judge and human experts, is constructed, which determines whether there exists a shift in perspective according to $N$-round debating context and results. Furthermore, we add both constraints on the number of total debating rounds and response tokens. Specifically, each team is assigned a total output token quota, with reminders of the remaining quota provided in each round, thereby creating a stressful evaluation environment. Since different models use different tokenizers, we instead compute the number of characters.
@@
Marbles. In the marbles scenario, the players first paired up and were each given ten marbles. The game’s objective was to gain the other player’s marbles through any game of their choice within the given time period. Inspired by this, we propose an asymmetric knowledge challenge, i.e., instructing the model to design questions that are most advantageous to oneself while targeting the opponent’s weaknesses. If the respondent answers incorrectly, the questioner gains one chip from the respondent and vice versa. The process ends when one side has lost all of its chips or the given number of rounds is reached. Note that the results of each round will be fed back to the questioner, enabling it to adjust its question-generation strategy. We employ a judging system similar to the tug of war game, where human experts are replaced with a string parsing algorithm.
@@
Glass Stepping Stones. The game involved contestants attempting to cross two parallel bridges by jumping across tempered glass panels while avoiding weaker panels of regular glass. Those who landed on a regular glass broke through the panel and fell to their elimination. Here, we transform this game to evaluate the proficiency of models in sequential decision-making under severe uncertainty. Specifically, the model is tasked with crossing a metaphorical bridge consisting of $N$ sequential steps. At each step $i\in\{1,\ldots,N\}$, there is a pair of glass panels, $\{{L_{i}},{R_{i}}\}$, with a safe path unknown to the model. Initially, each model determines its sequence of actions after learning the game rules. The first model $M_{1}$ to attempt the bridge acts as the information producer for the entire group and faces maximum uncertainty. Subsequent models $M_{k>1}$ act to effectively leverage the public information generated by their predecessors. If the number of survivors in each round is less than the total number of steps remaining required to pass, the surviving models will then re-select their movement order and resume the above process. The objective for each model $M_{k}$ is to derive an optimal policy $\pi_{k}^{*}$ that maximizes its survival probability given the actions of all preceding models:
@@
where $H_{k,i}$ denotes the public history available to model $M_{k}$ at $i$-th step. $k$ is the index for the model, representing its order in the sequence. This multi-agent design transforms a simple probability puzzle into a rich simulation of real-world competitive intelligence, where agents must learn not only from direct interaction with the environment but also from the successes and failures of others.
Figure 3: Survival rate of each model across six levels in Squid Game. We merely report the models that passed the first game.
The Final Squid Game. The ultimate showdown contains an attacker and a defender. The attacker wins by penetrating the defensive line to reach the target zone, whereas the defender secures victory by removing the attacker from the area of play. To this end, we elaborate a stress test as the final competition to evaluate the model’s robustness in safety alignment and its meta-cognitive abilities, which measures the model’s performance in a zero-sum, adversarial dynamic environment that mirrors the real-world interplay between LLM safety developers and malicious actors. To win the game, the attacker aims to analyze the cognitive process of the defender and identify logical loopholes, thereby crafting input jailbreaks to induce a violation. The model achieving the maximum number of successful attacks or defenses within a given round limit is declared the ultimate victor. Success in this game provides strong evidence of a model’s readiness for deployment in open-ended, adversarial real-world scenarios. See more details in App. B, and we provide prompt templates in App. C.2.
3 Experiments
3.1 Evaluation Setup
Participating Models. Our Squid Game includes 52 LLMs total, with a mix of top proprietary models and open-source models across various sizes. In particular, for proprietary models, we include OpenAI models such as GPT-5, GPT-4o, and o3 (OpenAI, 2024; 2025a; 2025b), Google models such as Gemini 2.5 Pro and Gemini 1.5 Flash (Mallick & Kilpatrick, 2025; Team et al., 2024), Anthropic models such as Claude 4.1 Opus and Claude 3.7 Sonnet (Anthropic, 2025a; c), xAI models such as Grok-3 and Grok-4 (x.AI, 2025a; b), and ByteDance models such as Seed 1.6 (ByteDance, 2025). For open-source models, we include models from Qwen (Qwen3-235B, Qwen2.5-{72B, 7B} (Yang et al., 2025a; Qwen, 2024)), DeepSeek (DS-R1, DS-V3 (Liu et al., 2024; Guo et al., 2025)), Llama3.x (Grattafiori et al., 2024), Kimi-K2 (Kimi et al., 2025), and GLM (GLM-{4, 4.5, 4.5-air} (GLM et al., 2024; Zeng et al., 2025)) families. App. C.1 provides a full list of models and citations.
Data and Implementation Details. We use default sampling settings, such as temperature and top_p, for all models. We allow for the maximum generation length possible for each model while instructing them to control their output length via the prompt in scenarios with resource constraints. The number of rounds or steps $N$ in all games, except for sugar honeycombs, is set to $20$. We conducted the whole Squid Game independently 20 times to mitigate randomness.
Table 1: The survival situation of each level in the Squid Game. $SR_{s}$ and $SR_{o}$ represent the stage survival rate and overall survival rate, respectively. We show the mean and standard deviation obtained from 20 independent Squid Games.
Game (#Avg./Max./Min. LLMs) Player Statistics $SR_{s}$ $SR_{o}$

### v2

@@ Abstract
The potential data contamination issue in contemporary large language models (LLMs) benchmarks presents a fundamental challenge to establishing trustworthy evaluation frameworks. Meanwhile, they predominantly assume benign, resource-rich settings, leaving the behavior of LLMs under pressure unexplored. In this paper, we introduce Squid Game, a dynamic and adversarial evaluation environment with resource-constrained and asymmetric information settings elaborated to evaluate LLMs through interactive gameplay against other LLM opponents. Squid Game consists of six elimination-style levels, focusing on multi-faceted abilities, including instruction-following, code, reasoning, planning, and safety alignment. We evaluate over 50 LLMs on Squid Game, presenting the largest behavioral evaluation study of general LLMs on dynamic adversarial scenarios. We observe a clear generational phase transition in performance in the same model lineage and find evidence that some models resort to speculative shortcuts to win the game, indicating the possibility of higher-level evaluation paradigm contamination in static benchmarks. We also compare prominent LLM benchmarks and Squid Game, highlighting that dynamic evaluation can serve as a complementary part for static evaluations. Project page: https://github.com/zijianchen98/LLM_Squid_Game.
Machine Learning, ICML
@@ 1 Introduction
Evaluation has always been an important cornerstone for large language model (LLM) development, with a proliferation of general benchmarks (Bai et al., 2024; Wang et al., 2024; White et al., 2025) and their extensions flourishing across multiple domains, covering specific disciplines (Sun et al., 2024; Chen et al., 2025a; Rein et al., 2024), reasoning (Valmeekam et al., 2023), code (Jain et al., 2025), safety (Zhang et al., 2024b), and multi-modality (Li et al., 2024a; Fu et al., 2024; Yue et al., 2024; Chen et al., 2025d, b) tasks. Despite these rapid advancements, evaluation methodologies have remained relatively stagnant. Current benchmarks predominantly conduct evaluations under benign environments with abundant computational resources (e.g., no quota limit) and stable interaction (e.g., fixed question-answer format), resulting in a pronounced gap between their high theoretical performance and practical utility. Moreover, most of them are static, closed-ended, and composed of knowledge-intensive tasks, making them susceptible to data contamination (Xu et al., 2024; Deng et al., 2024).
@@
Figure 1: A simplified tournament bracket of the Squid Game, showing the advancement path of competing models through successive elimination rounds. Squid Game comprises levels from static parallel to dynamic adversarial settings, allowing all-round evaluations.
@@
Motivated by these shortcomings, we introduce Squid Game, built on the following principles.
1) Elimination rather than Score. Traditional score-based benchmarks (Fu et al., 2024; Wang et al., 2024; Yue et al., 2024; Chen et al., 2025a) conduct evaluations by completing an examination independently in a static and resource-unlimited environment. Then, an absolute score is obtained to measure the model’s mastery over a specific body of knowledge, essentially answering the question “How much do you know?”. Such a scoring system could lead to the phenomenon of ranking fraud, where model developers may over-optimize for the specific question types, constituting a form of exam-oriented training. In Squid Game, we adopt a Battle Royale-style relative ranking system, where a certain percentage of the models may be eliminated in each round. Based on this, the difficulty of following challenges increases according to both the game settings and the survivors themselves. Here, a high ranking not only represents superior capabilities but also reflects a tactical advantage (e.g., fewer errors, better strategies) over specific opponents. Fig. 1 shows an example of dynamic evaluation results.
2) Resource Constraint. With the increasing scale and breadth of benchmarks, the demand for evaluation resources has also grown accordingly. Running thoroughly on MMMU-Pro (Yue et al., 2024) for a 72B LLM requires more than two GPUs, each with 80GB of VRAM, taking several dozen hours to complete. As for API models, evaluation costs can run into the tens of dollars, which is magnified by modern reasoning models that inflate costs by generating a large volume of verbose and often superfluous output tokens. To this end, we add the resource constraint (e.g., token or API call quota) in Squid Game to measure the utilization rate of tokens and inference efficiency. More importantly, this mechanism also exerts pressure on the model by providing live feedback on its diminishing resource pool.
3) Information Asymmetry. Mainstream benchmarks (Hendrycks et al., 2021; Cobbe et al., 2021; Yue et al., 2024; White et al., 2025) have almost all adopted multi-choice, binary judgment, or open-ended forms under fixed question-answer (QA) scenarios, where the core assumption is one of complete information, i.e., the prompt itself provides all necessary context. Furthermore, option-based evaluation is more susceptible to hallucination and output preferences (Chen et al., 2025d), resulting in biased observations. In contrast, we introduce an information asymmetry design that shifts the evaluative focus from “what a model knows” to “what it can do under uncertainty”. This further improves the depth of evaluation and can differentiate models with better reasoning, planning, and strategy generation capabilities.
4) Dynamic Adversarial Evaluation. Recent works (Chan et al., 2024; Gao et al., 2025) have demonstrated the feasibility of the interactive and collaborative evaluation paradigm for LLMs. Google’s recent chess tournament for LLMs (Google, 2025) serves as a good example. In Squid Game, we elaborate a series of levels (e.g., red-green light, tug of war, and marbles) that provide a holistic evaluation of all-round capabilities, ranging from benign instruction following to collaborative problem-solving and adversarial gaming in both offensive and defensive scenarios. This creates a self-evolving and never-saturating evaluation environment, where the difficulty automatically scales with the opponent’s intelligence, thus maintaining a constant challenge for cutting-edge models.
Figure 2: Overview of the six scenarios present in Squid Game.
With these principles in mind, we build Squid Game, a dynamic adversarial game-form benchmark that exposes LLMs to extreme stress-testing conditions with six evaluation scenarios.
Empirical Findings. We have evaluated 52 LLMs (28 proprietary and 24 open-source) across different Squid Game scenarios. Based on this study and the following analysis, we present novel empirical findings that have not been revealed in prior benchmarks.
@@
    Holistic Evaluation. We reveal that model dynamic performance is a multifaceted outcome, determined not only by task complexity but also by its architectural class (e.g., lightweight and inherent reasoning capabilities), rather than a monolithic scaling law.
@@
    Constructing Dynamic Evaluation. We observe a potential leakage in evaluation methodologies, which is as important as data contamination, highlighting the necessity of building dynamic evaluation frameworks.
@@
    Complementary to Static Evaluation. We observe relatively orthogonal results between Squid Game and static benchmarks, indicating the necessity of exploring the dynamic abilities of LLMs in the current landscape, since the performance on static benchmarks is approaching saturation.
2 Related Work
Data Contamination and Countermeasures. Benchmark data contamination is a critical issue encountered during the training and evaluation of LLMs, which can significantly impact the reliability of model’s performance scores, leading to untrustworthy, inflated results that do not accurately reflect its true capabilities (Magar & Schwartz, 2022; Xu et al., 2024; Deng et al., 2024). Early works (Radford et al., 2019; Brown et al., 2020) used high-order $n$-grams to detect overlapping content between the pre-training data and the evaluation datasets in GPT-2 and GPT-3. After analyzing 255 papers and considering OpenAI’s data usage policy, the authors in (Balloccu et al., 2024) report that the most prominently used LLMs today have been globally exposed to over 4.7M samples from 263 benchmarks. Recent work (Yang et al., 2025b) found a significant image overlap of over 84% and 33% between Seed-Bench (Li et al., 2024a) and pre-training datasets LAION-100M (Schuhmann et al., 2021) and CC3M (Sharma et al., 2018), respectively. In (Song et al., 2024), researchers expanded the concept of multimodal data contamination, as it pertains to the modality of data sources exposed to the MLLMs, and observed that proprietary models, such as Claude 3.5 Sonnet (Anthropic, 2024), show higher contamination levels than open-source models in ScienceQA (Lu et al., 2022) training and test sets. One promising solution to mitigating this problem is dynamic evaluation. DyVal (Zhu et al., 2024) leverages the structural advantage of directed acyclic graphs to dynamically generate evaluation samples with controllable complexities. Similarly, DARG (Zhang et al., 2024a) extracts the reasoning graphs of data points in current benchmarks and then perturbs the reasoning graphs to generate novel testing data. Additionally, various bootstrapping strategies (e.g., image editing and sentence rephrasing) (Yang et al., 2025b) with complexity control for both image and question modification are employed for dynamic evaluation. More recently, some studies (Wang et al., 2025; Chen et al., 2024) utilized multi-agent systems to generate real-time variable instances or reframe new ones for self-evolving benchmarks. However, these countermeasures predominantly focus on data reformation while neglecting the potential conceptual and methodological leakage. Rather than memorizing specific test instances, the model may have internalized the underlying solution patterns or heuristics for a given problem archetype from its training corpus. Our Squid Game aims to overcome these hurdles by introducing dynamic adversarial evaluation scenarios and neutralizing meta-knowledge leakage through information asymmetry and procedural generation.
@@ 3 Constructing the Squid Game
Games provide an intuitive and comparable signal of success. Unlike the unidirectional question-answer format-based evaluations, they compel models to demonstrate many capabilities, including strategic reasoning, planning, and survivability, under competitive scenarios, offering a stepwise visualization of their behavior (Hu et al., 2024; Zheng et al., 2025; Bailis et al., 2024; Olszewska & Risdal, 2025). Such game-based behavioral research is also quite popular in other fields. Previous prevailing Netflix series, Squid Game, a South Korean survival drama with critical acclaim, has been widely studied in the sociology domain (Aoun, 2022). Such high-pressure game settings serve to magnify each participant’s capabilities and human frailties, which satisfy the reform requirements of today’s LLM evaluation. In this work, we introduce Squid Game, a dynamic adversarial benchmark for general LLMs with six stressful evaluation scenarios, i.e., red-green light, sugar honeycombs, tug of war, marbles, glass stepping stones, and the final squid game, as shown in Fig. 2.
3.1 Game Design
@@
$L_{1}$: Red-Green Light. In this game, players can move on a green light but must freeze on a red light. Otherwise, they will be eliminated. Correspondingly, we focus on the accuracy and timeliness of LLMs’ instruction-following ability and adopt a turn-based, long-term task decomposition design. At the beginning of the game, the model receives a system prompt $P_{system}$ that defines the complete requirement of engagement, including the general description of the task and a rule for generating the security code. The game proceeds in turns. In each round, two types of commands are randomly generated with a certain probability $\rho$, namely “Continue” (green light) for generating the next chunk of the long task and “Interruption” (red light) for immediately stopping the main task and outputting the security code according to the predefined rule. A verification function is deployed to eliminate the models that do not meet the code requirements. The model then takes the output from the previous round, along with the next sub-task instruction, and proceeds.
Table 1: Characteristics of each level in Squid Game.
Level Elimin. Res. Cons. Info. Asy. Dyn. Adv. Win. Condition
$L_{1}$ Complete long task
$L_{2}$ Pass tests
$L_{3}$ Reverse argument
$L_{4}$ Question answer
$L_{5}$ Real-time observ.
$L_{6}$ Induce error
@@
$L_{2}$: Sugar Honeycombs. In this game, contestants were tasked with using a needle to carve a shape out of a piece of dalgona without cracking the honeycomb. Here, we focus on the precise control ability of models for performing delicate operations in demanding conditions. Specifically, we design a code sculpting scenario where the models are given a functionally complete but structurally suboptimal code $C_{in}$ with poor readability, high complexity, or redundant logic. This task is to refactor $C_{in}$ into a new version $C_{out}$ while satisfying a set of highly restrictive rules $\mathcal{R}$:
@@
where the survival function $\mathbb{I}_{survival}$ equals $1$ if the condition is met, and 0 if even a single test case fails. $\mathbb{I}(\cdot)$ denotes the indicator function and $C_{out}=M(C_{in},\mathcal{R})$. $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ includes the set of $n$ unit tests.
@@
$L_{3}$: Tug of War. This game assesses the teamwork capabilities across multiple models under random allocation. We involve two teams of three (i.e., the lead debater, supporting debater, and summarizer) debating for a given topic, where both sides engage in multiple rounds of language games, generating arguments, rebuttals, and attempting to steer the consensus of the central issue towards the position they had pre-determined (Estornell & Liu, 2024). A hybrid judging system, including a superior LLM judge (Zheng et al., 2023; Li et al., 2025) and human subjects, is built to determine whether there exists a shift in perspective according to $N$-round debating context and results. Furthermore, we add constraints on the number of total debating rounds and response tokens, where each team is assigned a fixed token quota, with reminders of the remaining quota provided in each round, thereby creating a stressful evaluation environment.
Figure 3: Survival rate of different LLMs across six levels in Squid Game. We merely report the models that passed the first game.
@@
$L_{4}$: Marbles. This game’s objective was to gain the other player’s marbles through any game of their choice within the given time period. Inspired by this, we propose an asymmetric knowledge challenge, i.e., instructing the model to design questions that are most advantageous to oneself while targeting the opponent’s weaknesses. If the respondent answers incorrectly, the questioner gains one chip from the respondent and vice versa. The process ends when one side has lost all of its chips or the given number of rounds is reached. Note that the results of each round will be fed back to the questioner, enabling it to adjust its question-generation strategy. We employ a similar judging system to the tug of war game.
@@
$L_{5}$: Glass Stepping Stones. In this game, we aim to evaluate the proficiency of models in sequential decision-making under severe uncertainty. Specifically, the model is tasked with crossing a metaphorical bridge consisting of $N$ sequential steps. At each step $i\in\{1,\ldots,N\}$, there is a pair of glass panels, $\{{L_{i}},{R_{i}}\}$, with a safe path unknown to the model. Initially, each model determines its sequence of actions after knowing the game rules. The first model $M_{1}$ to attempt the bridge acts as the information producer for the entire group and faces maximum uncertainty. Subsequent models $M_{k>1}$ act to effectively leverage the public information generated by their predecessors. If the number of survivors in each round is less than the total number of steps remaining required to pass, the surviving models will then re-select their movement order and resume the above process. The objective for each model $M_{k}$ is to derive an optimal policy $\pi_{k}^{*}$ that maximizes its survival probability given the actions of all preceding models:
@@
where $H_{k,i}$ denotes the public history available to model $M_{k}$ at $i$-th step. $k$ is the model’s order in the sequence. This multi-agent design transforms a simple probability puzzle into a simulation of real-world planning, where models must learn not only from direct environmental interaction but also from the successes and failures of others.
$L_{6}$: The Final Squid Game. The ultimate showdown contains an attacker and a defender, where we elaborate a stress test as the final competition to evaluate the model’s robustness in safety alignment and its meta-cognitive abilities. To win the game, the attacker aims to analyze the cognitive process of the defender and identify logical loopholes, thereby crafting input jailbreaks to induce a violation. The model achieving the maximum number of successful attacks or defenses within a given round limit is declared the ultimate victor. Success in this game may provide strong evidence of a model’s readiness for deployment in open-ended, adversarial real-world scenarios. Tab. 1 summarizes each level. See App. A for more details.
3.2 Dynamic Situational Broadcasting
To induce environmental stress and facilitate situational awareness, we further introduce dynamic situational broadcasting, which injects real-time tournament metadata into the prompt context. Let $\mathcal{M}=\{m_{1},m_{2},\dots,m_{k}\}$ be the set of participating models. At any time step $t$ within a game level, the global tournament state is defined as $\Omega_{t}=\langle\mathcal{N}_{t},\mathcal{E}_{t},\mathcal{L}_{t}\rangle$. Then, the input to a specific model $m_{i}\in\mathcal{N}_{t}$ is augmented as:
$\tilde{I}_{i,t}=\text{Concat}(I_{base},\Omega_{t},\mathcal{H}_{t-1})$ (3)
where $I_{base}$ denotes the task-specific instruction and $\mathcal{H}_{t-1}$ is the interaction history.
Table 2: The survival situation of each level in the Squid Game. “#Avg.”, “#Max.”, and “#Min.” denote the average, maximum, and minimum number of participating models, respectively. $SR_{s}$ and $SR_{o}$ represent the stage survival rate and overall survival rate, respectively. We show the mean and standard deviation obtained from 20 independent Squid Games.
Game #Avg. #Max. #Min. Player Statistics $SR_{s}$ $SR_{o}$